three times the larger of two numbers is equal to four times the smaller.the sum of the numbers is 21. please hhelp me find the numbers.
Answer by usyim88hk(158)

Let the small number be x and the bigger one be y
:
x+y = 21
4x = 3y
:
For this two equations, we can use substitution
:
y = 21 - x (change it into y-form)
3y = 4x (another y-form)
:
now we have to take y away so we have to have 3y on one equation and -3y on another equation to get 0y which is nothing (hard isn't it?)
:
-3y = -63 + 3x
3y = 4x
:
Now take away the ys by adding the equations
:
0 = 7x - 63
63 = 7x
x = 9
:
So now we know x and we can solve for y
:
3y = 4x
3y = 4(9)
3y = 36
y = 12
:
So the answers are 9 and 12